The mean is 15.2 and the standard deviation is 0.9. Find the probability that X is greater than 17.
Answer by stanbon(57347)

The mean is 15.2 and the standard deviation is 0.9. Find the probability that X is greater than 17.
-------
z(17) = (17-15.2)/0.9 = 2
------
P(x > 17) = P(z > 2) = normalcdf(2,100) = 0.0228
=====================================================
Cheers,
Stan H.